# Orbital magnetization in Sierpinski fractals

### v2

@@ Abstract
The more significant fractal-induced gaps and corresponding plateaus in orbital magnetization for both ZZ ST and AC ST are highlighted in blue shaded rectangles in Figure 2 (b) and Fig. 4 (c), respectively. Interestingly, for ST AC G(6), a large fractal gap overlaps with much of the 2D Haldane bulk gap, producing a wide plateau in the graph of orbital magnetization versus chemical potential. For both types of ST, the results obtained using the primary definition of orbital magnetization (Eq. 2), denoted by a continuous line, and those obtained from the local markers formulation (Eq. 3 and 8), denoted by circles, coincide. The local markers approach provides a transparent picture of the emergence of plateaus in orbital magnetization produced by fractal gaps. When the chemical potential $\mu$ passes through an energy range with no electronic states (fractal gap), the ground-state projector $\mathcal{P}$ remains invariant, and, since the whole-sample average of $\mathfrak{C}({\bf r})$ in a system with OBC $\int_{A}d{\bf r}\mathfrak{C}({\bf r})$] is enforced to vanish [21, 48], no change in orbital magnetization occurs. The analysis of the local Chern marker in distinct fractal-induced gaps for ST ZZ is presented in the supplementary material.
@@
These results illustrate the direct impact of the fractal geometry on the orbital magnetization. As the fractal generation increases, the orbital magnetization curve develops a more intricate structure, reflecting the emergence of a more complex hierarchy of energy gaps in the electronic spectrum.
@@
Final Remarks: The emergence of new fractal-induced gaps is a direct consequence of the ST geometry. Its iterative construction creates a self-organized hierarchy of boundaries, leading to quantum confinement that quantizes the electronic energy levels. The triangular geometry ensures each spatial side composed by a single, pure edge type either zigzag (ZZ) or armchair (AC), whose intrinsic properties dominate the electronic behavior. Consequently, switching the boundary type drastically modifies the energy spectra and orbital magnetization, which exhibit distinct plateaus within these fractal-induced gaps. In contrast to SCs fractals, which develop a dense, staircase-like spectrum and growing oscillations in orbital magnetization, the STs emergent gaps promote constant orbital magnetization plateaus. This shows how quantum confinement in fractal geometry can be used to explore novel properties of electronic orbital angular momentum, offering an interesting perspective in the emerging field of orbitronics [39]. Proposals and experimental realizations of fractal geometry are available for different types of physical systems [29, 50, 51, 8, 7, 52]. Furthermore, the Haldane model, used here as a prototypical example, can be realized in different contexts, such as magneto-optical lattices [53], ultracold fermion lattices [54], Fe-based insulators [55], and topoelectrical circuits [56], opening a perspective to study the phenomenology discussed here.
@@
Acknowledgments: The authors would like to thank the INCT de Nanomateriais de Carbono for providing support on the computational infrastructure. LLL thanks the CNPq scholarship. TPC acknowledges CNPq (Grant No. 305647/2024-5). AL thanks FAPERJ (Grant E-26/200.569/2023).

### v3

@@ V Discussion
The more significant fractal-induced gaps and corresponding plateaus in orbital magnetization for both ZZ ST and AC ST are highlighted in blue shaded rectangles in Fig. 2 (b) and (c), respectively. Interestingly, for ST AC G(4), a large fractal gap overlaps with much of the 2D Haldane bulk gap, producing a wide plateau in the graph of orbital magnetization versus chemical potential. For both types of ST, the results obtained using the primary definition of orbital magnetization (Eq. 2), denoted by a continuous line, and those obtained from the local markers formulation (Eq. 3 and 8), denoted by circles, coincide. The features of orbital magnetization in fractal systems cannot be understood from the usual k-space formulation, instead requiring a purely real-space perspective [58]. In particular, an elegant interpretation of the orbital-magnetization plateaus in ST fractals can be obtained from the local-marker formulation presented in Eqs. (38). When the chemical potential $\mu$ passes through an energy range with no electronic states (fractal induced gaps marked in blue on real-space electronic spectra), the ground-state projector $\mathcal{P}$ and its complement $\mathcal{Q}$ remains invariant. Consequently, $m_{\rm LC}({\bf r})=\frac{e}{\hbar c}\text{Im}\langle{\bf r}\big|\mathcal{P}x\mathcal{Q}\mathcal{H}\mathcal{Q}y\mathcal{P}\big|{\bf r}\rangle$, $m_{\rm IC}({\bf r})=-\frac{e}{\hbar c}\text{Im}\langle{\bf r}\big|\mathcal{Q}x\mathcal{P}\mathcal{H}\mathcal{P}y\mathcal{Q}\big|{\bf r}\rangle$, and their corresponding sample average, $M_{\rm IC}=A^{-1}\int_{A}m_{\rm IC}({\bf r})d{\bf r}$ and $M_{\rm LC}=A^{-1}\int_{A}m_{\rm LC}({\bf r})d{\bf r}$, remain unchanged by $\mu$ within this energy range. The only term in Eq. (3) that has an explicit dependence on $\mu$ is the term $m_{\rm BC}({\bf r})=\mu\frac{e}{2\pi\hbar c}\mathfrak{C}(\bf r)$. Although it is locally modified by the chemical potential, the average over the entire sample must vanish, $M_{\rm BC}=A^{-1}\int_{A}m_{\rm BC}({\bf r})d{\bf r}=0$. This follows from the fact that the average of quantity $\mathfrak{C}({\bf r})$ vanishes when integrated over the whole system $A^{-1}\int_{A}\mathfrak{C}({\bf r})d{\bf r}=0$ [see the analysis in Appendix A], a well-known property of the Chern marker introduced by Bianco and Resta [5, 58]. In Fig. 4 (a), we present the partial contribution of each term in Eq. (3) to the total orbital magnetization [Eq. (8)] for the illustrative case of AC ST G(4) fractal systems. As explained above, the $M_{\rm BC}$ contribution vanishes at any $\mu$. In the Sierpinski triangles, the $M_{\rm LC}$ and $M_{\rm IC}$ terms display plateaus that reflect the total orbital-magnetization plateaus reported in previous results. The same idea holds for the magnetization plateau in the other ST fractals.
Fig. 4 (b) and (c) also present the total orbital magnetization and its separated contributions for the illustrative cases of AC ST G(0) and SC G(3) systems. In these cases, total magnetization also arises from the competition between the itinerant ($M_{\text{IC}}$) and local ($M_{\text{LC}}$) contributions. The Berry-curvature contribution vanishes in all cases because the local Chern marker, when integrated over the whole sample, also vanishes.
@@
These results illustrate the direct impact of the fractal geometry on the orbital magnetization. As the fractal generation increases, the orbital magnetization curve develops a more intricate structure, reflecting the emergence of a more complex hierarchy of energy structures in the electronic spectrum.
VI Conclusion
@@
A remarkable feature of STs geometry is the emergence of fractal-induced gaps [45, 42]. Its iterative construction creates a self-organized hierarchy of boundaries, leading to quantum confinement that quantizes the electronic energy levels. Such gaps have been shown to lead to interesting consequences for topology [16] and the formation of spin polarization [45] in fractal systems. The orbital-magnetization plateaus found here add a new and noteworthy feature to these gaps. The triangular geometry ensures each spatial side composed by a single, pure edge type either zigzag (ZZ) or armchair (AC), whose intrinsic properties dominate the electronic behavior. Consequently, switching the boundary type drastically modifies the energy spectra and orbital magnetization, which exhibit distinct plateaus within these fractal-induced gaps. In contrast to SCs fractals, which develop a dense, staircase-like spectrum and growing fluctuations in orbital magnetization, the STs emergent gaps promote constant orbital magnetization plateaus. This shows how quantum confinement in fractal geometry can be used to explore novel properties of electronic orbital angular momentum, offering an interesting perspective in the emerging field of orbitronics [13]. Proposals and experimental realizations of fractal geometry are available for different types of physical systems [11, 22, 34, 28, 8, 40]. Furthermore, the Haldane model, used here as a prototypical example, can be realized in different contexts, such as magneto-optical lattices [1], ultracold fermion lattices [27], Fe-based insulators [29], and topoelectrical circuits [20], opening a perspective to study the phenomenology discussed here.
Acknowledgments
@@
The authors would like to thank the INCT de Nanomateriais de Carbono for providing support on the computational infrastructure. LLL thanks the CNPq scholarship. TPC acknowledges CNPq (Grant No. 305647/2024-5). AL thanks FAPERJ (Grant E-26/200.569/2023).
Appendix A Site-Resolved Local Chern Marker within Fractal-Induced Gaps
Differently from the $\bf k$}-space formula of orbital magnetization, valid for uniform bulk system, where plateaus are associated with a null Chern number [55], for the Sierpinski triangles the plateaus in orbital magnetization can coexist with a non-null Local Chern Marker as can be understood with a real space projection. Fig. A1 shows the site-resolved local Chern marker (LCM) calculated for ST considering distinct fractal generations and chemical potentials. In panel (a) of Fig. A1, we show ZZ ST G(0) at half filling. In this case, the center of the sample has a quantized positive value of the Chern marker, whereas at the edges, the Chern marker assumes negative values. This profile is expected for the local Chern marker for a uniform system subjected to OBC in the Chern insulator phase. The negative values of the marker at the edges compensate for the positive value in the middle of the sample in such a way that the whole-sample average of the marker vanishes, as discussed in Ref. [5]. In panels (b), (c), and (d) of Fig. A1, we show the site-resolved local markers for higher generations of ZZ ST with the chemical potential in the middle of the fractal induced gaps, i.e., (b) ZZ ST G(3) with $\mu=-0.07t_{1}$, (c) ZZ ST G(4) with $\mu=+0.18t_{1}$, and (d) ZZ ST G(3) with $\mu=+0.18t_{1}$. For all three cases, the local Chern marker assumes finite values at the sites, but the whole-sample average of the marker also vanishes.
Figure A1: Site-resolved local Chern marker for (a) ZZ ST G(0) at half filing, (b) ZZ ST G(3) with $\mu=-0.07t_{1}$, (c) ZZ ST G(3) with $\mu=+0.18t_{1}$ and (d) ZZ ST G(4) with $\mu=+0.18t_{1}$. Note that both negative and positive values, marked as black and purple sites, respectively, are present in the site-resolved map. The legend color bar is cut at the range, which is between 0 and 1 values. The negative values of the marker compensate for the positive values in such a way that the whole-sample average of $\mathfrak{C}({\bf r})$ vanishes.
Appendix B Additional results on orbital magnetization in ST
Figure B1: Energy spectrum and orbital magnetization for model parameters set at point (2) in the trivial phase of the topological phase diagram shown in Fig. 1(a): $t_{2}=t_{1}/3$, $V_{ab}/t_{2}=8.0$ and $\phi=0.5\pi$. The results are shown for AC ST (a) G(0) and (b) G(4). The Haldane bulk gap is highlighted in yellow, and the fractal induced gaps in blue.
Here we extend our analysis by exploring a trivial point in the bulk Haldane phase diagram shown in Fig. 1 (a). In the 2D (bulk) system, the Haldane model presents particle-hole symmetry for $\phi=0.5\pi$. Then we consider this value of $\phi$ and a representative value of the sublattice potential $V_{ab}/t_{2}=8.0$, which drives the system to the topologically trivial phase, marked as point (2) in Fig. 1 (a). In Fig. B1 we present the electronic spectra and the orbital magnetization as a function of $\mu$ for the AC ST G(0) and G(4) in panels (a) and (b), respectively. The overall phenomenology follows the one detailed in the main text, except that in the trivial-insulator case, the states invading the Haldane bulk-gap region (yellow area) are absent in the finite-system energy spectra. In addition, for the parameters shown in Fig. B1, fractality slightly enlarges the central gap. In the bulk-state energy bands, fractal-induced gaps characterized by orbital magnetization plateaus also occur, being represented in the blue regions. A similar behavior is also observed in ZZ ST in the topologically trivial phase and is therefore not presented here to avoid repetition.
Figure B2: Energy spectrum and orbital magnetization of (a) AC and (b) ZZ ST G(4) lattices as a function of $\phi$. In both cases $t_{2}=1/3$ and $V_{ab}/t_{2}=3.0$. The vertical red lines in the plots of the spectra indicate the value $\phi=\pi/2$. In the plots of magnetization normalized by the area $A_{c}$, three distinct electronic occupations are considered: $\nu=0.50$ (triangles), $0.33$ (circles) and $0.20$ (squares). The color scale indicates the Fermi level $E_{F}$ (in units of $t_{1}$) for each occupation. The shaded gray areas designate the Chern insulator regions in the 2D Haldane phase diagram. The remaining portions in white corresponds to a trivial insulator.
In Fig. B2 we present the evolution of the electronic spectra and the orbital magnetization at fixed electronic occupations $\nu$ (0.20,0.33,0.50) as the phase $\phi$ of the Haldane model is varied. The results are shown for two illustrative cases, the AC ST G(4) and ZZ ST G(4) with $t_{2}=t_{1}/3$ and $V_{ab}/t_{2}=3.0$. Notably, for triangular systems with AC edges, the number of sites on sublattices A and B is equal. Nevertheless, in triangular systems with ZZ edges, there is an excess of sites on one of the sublattices. For this reason, the electronic spectra of AC ST G(4) preserve particle–hole symmetry when $\phi=\pi/2$, whereas this symmetry is lost for ZZ ST G(4) with finite $V_{ab}$. This effect arising from an imbalance in the number of sites in different sublattices and its signatures in the electronic spectra of fractals are discussed in detail in Ref. [42]. This distinction between fractals with different terminations is reflected in the contrasting magnetization curves as a function of $\phi$ for similar electronic occupations, as reported in Fig. B2.